# SegmentDreamer: Towards High-fidelity Text-to-3D Synthesis with Segmented Consistency Trajectory Distillation

### v1

@@ 7 Conclusion
Potential Negative Impact. SegmentDreamer focuses on content systhesis, which may be used to create misleading content.
@@
To provide proofs of SegmenDreamer and visual comparisons of state-of-the-arts, this supplementary material includes the following contents:
@@
  • Section C: How to Connect SCTD with SDS
  • Section D: Computation Reduction Trick
  • Section E: Proof of Upper Bound of Distillation Error
  • Section F: Visual comparisons of State-of-the-art
  • Section G: Limitation and Potential Negative Impacts
@@ B Flaws in the Consistency Function of Guided Consistency Sampling
Clearly, without the target timestep in $\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t,\mathbf{y})$, optimizing a 3D representation $\boldsymbol{\theta}$ to satisfy both conditions in Eq. 16 for all intervals $[e,t]$ is invalid. For GCS, as the number of training steps increases, the above unreasonable phenomena occur frequently, potentially resulting in poor distillation results.
C How to Connect SCTD with SDS?
@@
$\hat{\boldsymbol{\epsilon}}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t,\mathbf{y})-\boldsymbol{\epsilon}=\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\mathbf{y})-\boldsymbol{\epsilon}-\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})+\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})$ (18)
$+\omega(\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\mathbf{y})-\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\emptyset)),$
@@
D Computation Reduction Trick
@@
Based on Eq. 21, we can readily derive Eq. 9 by following the procedure described in App. C.
E Proof of Upper Bound of Distillation Error
@@
where $\tilde{\mathbf{z}}_{s}^{\boldsymbol{\Phi}}=\boldsymbol{\Phi}(\mathbf{z}_{s_{m}}^{\mathbf{c}},s_{m},s,\mathbf{y})$. According to the condition, we have
@@
$=\hat{\mathbf{z}}_{s_{m}}^{\boldsymbol{\Phi}}-\tilde{\mathbf{z}}_{s_{m}}^{\boldsymbol{\Phi}}=\mathbf{z}_{s_{m}}^{\mathbf{c}}-\mathbf{z}_{s_{m}}^{\mathbf{c}}=0,$
@@
F Visual comparisons of State-of-the-arts
We also provide visual comparisons of State-of-the-arts, as shown in Fig. 8. We provide additional qualitative comparisons with DreamFusion [31], LucidDreamer [23], Consistent3D [48], and ConnectCD [22]. CFG scales are set to 100, 7.5, 20$\sim$40, 7.5, 7.5, respectively. Our approach yields results with high quality within the shortest time. In Table 2, we provide the detailed prompts of examples.
G Limitation and Potential Negative Impact
Limitation. SegmentDreamer is primarily designed for single-instance generation and performs suboptimally in multi-instance generation.
Potential Negative Impact. SegmentDreamer focuses on content synthesis, which could potentially be misused to create misleading or deceptive content.
Figure 8: Additional qualitative comparisons with DreamFusion [31], LucidDreamer [23], Consistent3D [48], and ConnectCD [22]. CFG scales are set to 100, 7.5, 20$\sim$40, 7.5, 7.5, respectively. Our approach yields results with high quality within the shortest time. Please zoom in for details.
Table 2: Prompt of examples.
Prompt
1 A cat with a mullet
2 A pig wearing a backpack
3 A DSLR photo of an origami crane
4 A photo of a mouse playing the tuba
5 An orange road bike
6 A ripe strawberry
7 A DSLR photo of the Imperial State Crown of England
8 A photo of a wizard raccoon casting a spell
9 A DSLR photo of a corgi wearing a top hat
10 A rabbit, animated movie character, high-detail 3D model
11 A panda rowing a boat
12 A highly detailed sand castle
13 A DSLR photo of a chimpanzee dressed like Henry VIII king of England
14 A photo of a skiing penguin wearing a puffy jacket, highly realistic DSLR photo
15 A blue poison-dart frog sitting on a water lily
16 A DSLR photo of a bear dressed in medieval armor
17 A DSLR photo of a squirrel dressed like a clown
18 A plush toy of a corgi nurse
19 A humanoid robot playing the violin
20 A DSLR photo of a bear dressed as a lumberjack
21 A blue motorcycle
22 Michelangelo style statue of an astronaut
23 A DSLR photo of a chow chow puppy
24 A DSLR photo of cats wearing eyeglasses
25 A red panda
26 A DSLR photo of an elephant skull
27 An amigurumi bulldozer
28 A typewriter
29 A red-eyed tree frog, low poly
30 A DSLR photo of a chimpanzee wearing headphones
31 A robot made out of vegetables
32 A DSLR photo of a red rotary telephone
33 A DSLR photo of a blue lobster
34 A DSLR photo of a squirrel flying a biplane
35 A DSLR photo of a baby dragon hatching out of a stone egg
36 A DSLR photo of a bear dancing ballet
37 A plate of delicious tacos
38 A DSLR photo of a car made out of cheese
39 A yellow school bus
40 A DSLR photo of a shiny beetle

### v2

@@ 7 Conclusion
Potential Negative Impact. SegmentDreamer focuses on content synthesis, which may be used to create misleading content.
Acknowledgements
This project is supported by the National Natural Science Foundation of China (12326618, U22A2095), the National Key Research and Development Program of China (2024YFA1011900), the Major Key Project of PCL (PCL2024A06), and the Project of Guangdong Provincial Key Laboratory of Information Security Technology (2023B1212060026).
@@ Acknowledgements
To provide proofs of SegmentDreamer and visual comparisons of state-of-the-arts, this supplementary material includes the following contents:
@@
  • Figure 1: Oversaturation and Artifacts in GCS
  • Section D: How to Connect SCTD with SDS
  • Section E: Computation Reduction Trick
  • Section F: Proof of Upper Bound of Distillation Error
  • Section G: Visual Comparisons of State-of-the-arts
@@ B Flaws in the Consistency Function of Guided Consistency Sampling
[…]
C Oversaturation and Artifacts in GCS
Figure 1: Analysis on self- and cross-consistency in GCS (a), and visual analysis of oversaturation and artifacts (b)-(d).
Fig. 1 shows that the cross-consistency loss $\mathcal{L}_{\text{CG}}+\mathcal{L}_{\text{CP}}$ (with $\mathcal{L}_{\text{CP}}$ dominating) greatly exceeds the self-consistency loss $\mathcal{L}_{\text{CC}}$. As shown in Fig. 1, we scale down $\mathcal{L}_{\text{CP}}$ (c) and observe that the oversaturation (b) is alleviated, but undesirable geometry persists (red circle). This suggests that such oversaturation is brought from an excessively large value of cross-consistency loss, i.e., the “excessive conditional guidance,” which has been analyzed in Sec. 4.1. In contrast, our $\mathcal{L}_{\text{SCTD}}$ (d) successfully addresses the above issues.
D How to Connect SCTD with SDS?
@@
$\hat{\boldsymbol{\epsilon}}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t,\mathbf{y})-\boldsymbol{\epsilon}=\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\mathbf{y})-\boldsymbol{\epsilon}-\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})+$ (5)
$\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})+\omega(\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\mathbf{y})-\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\emptyset)),$
@@
E Computation Reduction Trick
@@
Based on Eq. 8, we can readily derive Eq. 9 by following the procedure described in App. D.
F Proof of Upper Bound of Distillation Error
@@
where $\tilde{\mathbf{z}}_{s}^{\boldsymbol{\Phi}}=\boldsymbol{\Phi}(\mathbf{z}_{s_{m}},s_{m},s,\mathbf{y})$. According to the condition, we have
@@
$=\hat{\mathbf{z}}_{s_{m}}^{\boldsymbol{\Phi}}-\tilde{\mathbf{z}}_{s_{m}}^{\boldsymbol{\Phi}}=\mathbf{z}_{s_{m}}-\mathbf{z}_{s_{m}}=0,$
@@
G Visual Comparisons of State-of-the-arts
We also present additional visual comparisons with state-of-the-art methods, as shown in Fig. 2 and Fig. 3. We provide additional qualitative comparisons against DreamFusion [31], LucidDreamer [23], Consistent3D [48], ConnectCD [22], Magic3D [24], Fantasia3D [3], and CSD [53]. As shown, our method clearly outperforms others visually.
Figure 2: Additional qualitative comparisons with DreamFusion [31], LucidDreamer [23], Consistent3D [48], and ConnectCD [22]. CFG scales are set to 100, 7.5, 20$\sim$40, 7.5, 7.5, respectively. Our approach yields results with high quality. Please zoom in for details.
Figure 3: Additional qualitative comparisons with Magic3D [31], Fantasia3D [23], and CSD [48]. Our approach yields results with high quality. Please zoom in for details.
Table 1: 40 prompts for evaluation.
Column 1 Column 2
1. A cat with a mullet 21. A blue motorcycle
2. A pig wearing a backpack 22. Michelangelo style statue of an astronaut
3. A DSLR photo of an origami crane 23. A DSLR photo of a chow chow puppy
4. A photo of a mouse playing the tuba 24. A DSLR photo of cats wearing eyeglasses
5. An orange road bike 25. A red panda
6. A ripe strawberry 26. A DSLR photo of an elephant skull
7. A DSLR photo of the Imperial State Crown of England 27. An amigurumi bulldozer
8. A photo of a wizard raccoon casting a spell 28. A typewriter
9. A DSLR photo of a corgi wearing a top hat 29. A red-eyed tree frog, low poly
10. A rabbit, animated movie character, high-detail 3D model 30. A DSLR photo of a chimpanzee wearing headphones
11. A panda rowing a boat 31. A robot made out of vegetables
12. A highly detailed sand castle 32. A DSLR photo of a red rotary telephone
13. A DSLR photo of a chimpanzee dressed like Henry VIII king of England 33. A DSLR photo of a blue lobster
14. A photo of a skiing penguin wearing a puffy jacket, highly realistic DSLR photo 34. A DSLR photo of a squirrel flying a biplane
15. A blue poison-dart frog sitting on a water lily 35. A DSLR photo of a baby dragon hatching out of a stone egg
16. A DSLR photo of a bear dressed in medieval armor 36. A DSLR photo of a bear dancing ballet
17. A DSLR photo of a squirrel dressed like a clown 37. A plate of delicious tacos
18. A plush toy of a corgi nurse 38. A DSLR photo of a car made out of cheese
19. A humanoid robot playing the violin 39. A yellow school bus
20. A DSLR photo of a bear dressed as a lumberjack 40. A DSLR photo of a shiny beetle